OpenStudy (anonymous):

Determine if the following function is even, odd, or neither. f(x) = –9x^4 + 5x + 3 show me steps please.

OpenStudy (anonymous):

you got an even power, namely 4, and an odd power, namely 1, so neither

OpenStudy (anonymous):

your teacher want you to do this \[f(-x)=-9(-x)^4+5(-x)+3=-9x^4-5x+3\] and since \[f(-x)\neq f(x)\] it is not even also since \(f(-x)\neq -f(x)\) it is not odd

OpenStudy (anonymous):

notice that \(-f(x)=-(-9x^4+5x+3)=9x^4-5x-3\) which is not the same as \(f(-x)=-9x^4-5x+3\)

OpenStudy (anonymous):

what do you mean f(−x)≠f(x)

OpenStudy (anonymous):

\(f(x)=-9x^4+5x+3\) right? and we computed \(f(-x)=-9x^4-5x+3\)

OpenStudy (anonymous):

so they are not equal because one has \(5x\) and the other has \(-5x\)

OpenStudy (anonymous):

because they are not the same, that means \(f\) is not even

OpenStudy (anonymous):

here is an example of an even function: \[g(x)=x^4+x^2+1\] in this example \[g(-x)=(-x)^4+(-x)^2+1=x^4+x^2+1=g(x)\] so that one is even

OpenStudy (anonymous):

oh I see and how do we know it's not odd